# Efficient Prior Selection in Gaussian Process Bandits with Thompson Sampling

### v1

@@ Abstract
Gaussian process (GP) bandits provide a powerful framework for solving blackbox optimization of unknown functions. The characteristics of the unknown function depends heavily on the assumed GP prior. Most work in the literature assume that this prior is known but in practice this seldom holds. Instead, practitioners often rely on maximum likelihood estimation to select the hyperparameters of the prior - which lacks theoretical guarantees. In this work, we propose two algorithms for joint prior selection and regret minimization in GP bandits based on GP Thompson sampling (GP-TS): Prior-Elimination GP-TS (PE-GP-TS) and HyperPrior GP-TS (HP-GP-TS). We theoretically analyze the algorithms and establish upper bounds for their respective regret. In addition, we demonstrate the effectiveness of our algorithms compared to the alternatives through experiments with synthetic and real-world data.
@@ 1 Introduction
In Gaussian process bandits, we consider a variant of the multi-armed bandit problem where the arms are correlated and their expected reward is sampled from a Gaussian process (GP). The flexibility of GPs have made GP bandits applicable in a wide range of areas that need to optimize blackbox functions with noisy estimates, including machine learning hyperparameter tuning Turner et al. (2021), drug discovery Hernández-Lobato et al. (2017); Pyzer-Knapp (2018), chemical design Griffiths & Miguel Hernández-Lobato (2020), battery charging protocols Jiang et al. (2022), online advertising Nuara et al. (2018), and portfolio optimization Gonzalvez et al. (2019). However, most of the theoretical results in the literature assume that the GP prior is known but this is seldom the case in practical applications. Even with expert domain knowledge, selecting the exact prior to use can be a difficult task. Most practitioners tend to utilize maximum likelihood estimation (MLE) to identify suitable prior parameters. However, in a sequential decision making problem MLE is not guaranteed to recover the correct parameters.
@@
In the literature, Wang & de Freitas (2014); Berkenkamp et al. (2019); Ziomek et al. (2024a) provided algorithms with theoretical guarantees when the kernel lengthscale is unknown. More recently, Ziomek et al. (2024b) introduced an elimination-based algorithm with theoretical guarantees for an arbitrary set of discrete priors. Their algorithm, Prior-Elimination GP-UCB (PE-GP-UCB), selects the arm and prior which provide the most optimistic upper confidence bound (UCB). If a prior generates too many incorrect predictions, then it may be eliminated. The previous work has focused on optimistic UCB methods which are known to over-explore.
@@
In this work, we investigate the use of Thompson sampling for solving GP-bandit problems with unknown priors and we propose two algorithms. The first algorithm, Prior-Elimination GP-TS (PE-GP-TS), is an extension of PE-GP-UCB that replaces the doubly optimistic selection rule with posterior sampling and one less layer of optimism. We obtain a regret bound for PE-GP-TS of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ where $T$ is the horizon and $\hat{\gamma}_{T}$ is the worst-case maximum information gain, which matches that of PE-GP-UCB. The second algorithm, HyperPrior GP-TS (HP-GP-TS), uses bi-level posterior sampling to efficiently explore the priors and arms. We derive a regret bound of order ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma}_{T}})$ where $\bar{\gamma}_{T}$ is the maximum information gain averaged across a hyperprior. The obtained bound matches that of standard GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ whenever the hyperprior is deterministic. Finally, we evaluate our methods on three sets of synthetic experiments and two experiments with real-world data. Across the experiments, our Thompson sampling based methods outperform PE-GP-UCB. Additionally, we analyze the priors selected by the algorithms and observe that HP-GP-TS selects the correct prior more often than the other algorithms.
@@
  • We provide theoretical regret bounds of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma_{T}}})$ and ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma_{T}}})$ for PE- and HP-GP-TS, respectively.
  • We experimentally evaluate our algorithms on both synthetic and real-world data, demonstrating that they achieve superior performance by avoiding the optimistic exploration of PE-GP-UCB.
@@ 2 Background and problem statement
Figure 1: Agent deciding between priors based on observed data. The unknown function $f$ is sampled from a periodic prior and the observed data is in the top left. The agent discards the priors on the right since they do not match the periodic observations.
@@
We consider a sequential decision making problem where an agent repeatedly selects among a set of arms and receives a random reward whose mean depends on the selected arm and is unknown to the agent. The goal of the agent is to maximize the cumulative sum of rewards over a finite time horizon. We assume that the distribution of the means, the prior, is sampled from a set of priors, the hyperprior. An effective agent must distinguish which prior the means are sampled from to ensure it explores efficiently. We illustrate this in Fig. 1.
Now, let us formally state the problem. Let ${\mathcal{X}}\subseteq[0,r]^{d}\subset{\mathbb{R}}^{d}$ denote the finite set of arms and $P$ a finite set of priors with associated prior mean and kernel functions $\mu_{1,p}:{\mathcal{X}}\mapsto{\mathbb{R}}$ and $k_{1,p}:{\mathcal{X}}\times{\mathcal{X}}\mapsto{\mathbb{R}}$, $\forall p\in P$. Let $p^{*}\in P$ denote the true prior and assume the expected reward function $f:{\mathcal{X}}\mapsto{\mathbb{R}}\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ is a sample from a Gaussian process with prior $p^{*}$. Both the function $f$ and the true prior $p^{*}$ are considered unknown. We will consider two settings: In the frequentist selection setting, the prior $p^{*}\in P$ is picked arbitrarily. In the Bayesian selection setting, the prior is sampled from a hyperprior $p^{*}\sim\mathcal{P}(P)$. To simplify notation, let $P_{0}$ denote the hyperprior.
Let $T$ denote the horizon. For time step $t=1,2,\ldots,T$, the agent selects an arm $x_{t}\in{\mathcal{X}}$ and observes the reward $r_{t}=f(x_{t})+\epsilon_{t}$ where $\{\epsilon_{t}\}_{t=1}^{T}$ are i.i.d. zero-mean Gaussian noise with variance $\sigma^{2}$. The goal of the agent is to select a sequence of arms $\{x_{t}\}_{t=1}^{T}$ that minimizes the regret $R(T)=\sum_{t\in[T]}f(x^{*})-f(x_{t})$ where $[T]=\{1,\ldots,T\}$ and $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}f(x)$. In the Bayesian selection setting, we evaluate the agent based on the Bayesian regret ${\text{BR}}(T)={\mathbb{E}}\left[R(T)\right]$ where the expectation is taken over the prior $p^{*}$, the expected reward function $f$, the noise $\{\epsilon_{t}\}_{t=1}^{T}$ and the potentially stochastic selection of arms.
@@
A Gaussian process $f(x)\sim{\mathcal{GP}}(\mu,k)$ is a collection of random variables such that for any subset $\{x_{1},\ldots,x_{n}\}\subset{\mathcal{X}}$, the vector $[f(x_{1}),\ldots f(x_{n})]\in{\mathbb{R}}^{n}$ has a multivariate Gaussian distribution. The probabilistic nature of GPs makes them very useful for defining and solving bandit problems where the arms are correlated. Given $t$ observations $\{(x_{i},y_{i})\}_{i=1}^{t}$, the posterior mean and kernel functions of a Gaussian process ${\mathcal{GP}}(\mu,k)$ are given by
@@
$k_{t}(x,\tilde{x})$ $=k(x,\tilde{x})-{\mathbf{k}}^{\top}\left({\mathbf{K}}+\sigma^{2}I\right)^{-1}\tilde{{\mathbf{k}}}.$ (2)
Above, ${\mathbf{k}},\tilde{\mathbf{k}}\in{\mathbb{R}}^{t}$ are vectors such that $({\mathbf{k}})_{i}=k(x_{i},x)$ and $(\tilde{{\mathbf{k}}})_{i}=k(x_{i},\tilde{x})$. Additionally, ${\mathbf{y}},{\bm{\mu}}\in{\mathbb{R}}^{t}$ are also vectors such that $({\mathbf{y}})_{i}=y_{i}$ and $({\bm{\mu}})_{i}=\mu(x_{i})$. The gram matrix is denoted by ${\mathbf{K}}\in{\mathbb{R}}^{t\times t}$ where $({\mathbf{K}})_{i,j}=k(x_{i},x_{j})$. Let $\mu_{t,p}$ and $k_{t,p}$ denote the posterior mean and kernel for a Gaussian process with prior $p\in P$ at time $t$ and let $\sigma^{2}_{t,p}(x)=k_{t,p}(x,x)$ denote the posterior variance at time $t$.
The kernel $k$ determines important characteristics of the functions $f$ and in the following we provide some examples. The RBF kernel, $k(x,\tilde{x})=\exp(-||x-\tilde{x}||^{2}/\ell^{2})$ guarantees that $f$ is smooth. The lengthscale parameter $\ell>0$ determines how quickly $f$ changes, smaller values lead to more fluctuations. The rational quadratic (RQ) kernel $k(x,\tilde{x})=\left(1+\frac{||x-\tilde{x}||^{2}}{2\alpha\ell^{2}}\right)^{-\alpha}$ where $\alpha>0$ is a mixture of RBF kernels with varying lengthscales. The Matérn kernel Matérn (1986) $k(x,\tilde{x})=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}||x-\tilde{x}||}{\ell}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}||x-\tilde{x}||}{\ell}\right)$ where $\nu>0$ is the smoothness parameter that imposes that $f$ is k-times differentiable if $\nu>k$ for integer $k$. The functions $\Gamma(\nu)$ and $K_{v}$ correspond to the gamma function and a modified Bessel function Williams & Rasmussen (2006). The periodic kernel $k(x,\tilde{x})=\exp\left(-\frac{1}{2}\sum_{i=1}^{d}\sin^{2}(\frac{\pi}{\rho}(x-\tilde{x}))/\ell\right)$ generates smooth and periodic functions with period $\rho>0$ Mackay (1998). The linear kernel $k(x,\tilde{x})=vx^{\top}\tilde{x}/\nu$ generates linear functions where $v$ is the variance parameter.
@@
The maximal information gain (MIG) is a measure of reduction in uncertainty of $f$ after observing the most informative data points up to a specified size. The MIG commonly occurs in regret bounds for GP bandit algorithms Srinivas et al. (2012); Vakili et al. (2021) and its growth rate is strongly determined by the prior kernel of the GP. Hence, we will define the MIG for any fixed GP prior $p\in P$. Let ${\mathbf{y}}_{A}$ denote noisy observations of $f$ at the locations $A\subset{\mathcal{X}}$. Then, the MIG given prior $p\in P$, $\gamma_{T,p}$, is defined as
@@
where $I_{p}({\mathbf{y}}_{A};f)=H({\mathbf{y}}_{A}|p)-H({\mathbf{y}}_{A}|f,p)$ is the mutual information between ${\mathbf{y}}_{A}$ and $f$ given $p$, and $H(\cdot)$ denotes the entropy. To aid our analysis later, we also define the worst-case MIG as $\hat{\gamma}_{T}:=\max_{p\in P}\gamma_{T,p}$ and the average MIG as $\bar{\gamma}_{T}:={\mathbb{E}}_{p\sim P_{0}}[\gamma_{T,p}]$. For the RBF and Matérn kernels, $\gamma_{T,p}={\mathcal{O}}(\log^{d+1}(T))$ and $\gamma_{T,p}={\mathcal{O}}(T^{\frac{d}{2\nu+d}}\log^{\frac{2\nu}{2\nu+d}}(T))$ Srinivas et al. (2012); Vakili et al. (2021).
@@
Wang & de Freitas (2014) was the first work that presented regret bounds for Gaussian process bandits with unknown lengthscale. Wang & de Freitas studied the popular Expected Improvement algorithm Močkus (1975). But their algorithm requires a lower bound on the lengthscale and their regret bound depends on the worst-case MIG. Later work by Berkenkamp et al. (2019) introduced Adaptive GP-UCB (A-GP-UCB) that continually lowers lengthscale parameter. Once the lengthscale parameter is sufficently small, then the function $f$ will lie within the reproducing kernel Hilbert space (RKHS) and then the regular GP-UCB theory can be applied. However, A-GP-UCB lacks a stopping mechanism and will overexplore as the lengthscale continues to shrink. Recent work by Ziomek et al. (2024b) introduced Prior-Elimination GP-UCB (PE-GP-UCB) for time-varying GP-bandits with unknown prior. Unlike the work before, the regret bound of PE-GP-UCB holds for arbitrary types of hyperparametersin the GP prior. PE-GP-UCB is doubly optimistic and selects the prior and arm with the highest upper confidence bound. PE-GP-UCB tracks the cumulative prediction error made by the selected priors and eliminates priors that exceed a threshold level.
Other works have introduced regret balancing algorithms that maintain a set of base learning algorithms and balance their selection frequency to achieve close to optimal regret Abbasi-Yadkori et al. (2020); Pacchiano et al. (2020). Ziomek et al. (2024a) built on this idea and introduced length-scale balancing GP-UCB which can adaptively explore smaller lengthscales but can return to longer ones, unlike A-GP-UCB.
Wang et al. (2018) provided regret bounds for probability of improvement Kushner (1964) and GP-UCB when the GP prior is learnt from offline data, which they called Meta BO. The aforementioned works are based on UCB, EI, PI or regret balancing. However, another line of work has studied Thompson sampling in standard and linear bandits with unknown prior distribution Kveton et al. (2021); Basu et al. (2021); Hong et al. (2022); Li et al. (2024). In their setting (meta or hierarchical bandits), the agent plays multiple bandit instances, either simultaneously or sequentially. The unknown means are sampled from the same (unknown) prior and by gathering knowledge across instances, the agent can solve later instances more efficiently once it has identified the prior. We emphasize that these methods have been studied only for standard stochastic and linear bandits, not for GP bandits.

### v2

@@ Abstract
Gaussian process (GP) bandits provide a powerful framework for performing blackbox optimization of unknown functions. The characteristics of the unknown function depend heavily on the assumed GP prior. Most work in the literature assume that this prior is known but in practice this seldom holds. Instead, practitioners often rely on maximum likelihood estimation to select the hyperparameters of the prior - which lacks theoretical guarantees. In this work, we propose two algorithms for joint prior selection and regret minimization in GP bandits based on GP Thompson sampling (GP-TS): Prior-Elimination GP-TS (PE-GP-TS) and HyperPrior GP-TS (HP-GP-TS). We theoretically analyze the algorithms and establish upper bounds for the regret of HP-GP-TS. In addition, we demonstrate the effectiveness of our algorithms compared to the alternatives through experiments with synthetic and real-world data.
@@ 1 Introduction
In Gaussian process bandits, we consider a variant of the multi-armed bandit problem where the arms are correlated and their expected reward is sampled from a Gaussian process (GP). The flexibility of GPs have made GP bandits applicable in a wide range of areas that need to optimize blackbox functions with noisy estimates, including machine learning hyperparameter tuning (Turner et al., 2021), drug discovery (Hernández-Lobato et al., 2017; Pyzer-Knapp, 2018), online advertising (Nuara et al., 2018), portfolio optimization (Gonzalvez et al., 2019) and energy-efficient navigation (Sandberg et al., 2025). Most of the theoretical results in the literature assume that the GP prior is known but this is seldom the case in practical applications. Even with expert domain knowledge, selecting the exact prior to use can be a difficult task. Most practitioners tend to utilize maximum likelihood estimation (MLE) to identify suitable prior parameters. However, in a sequential decision making problem MLE is not guaranteed to recover the correct parameters.
@@
In the literature, Wang & de Freitas (2014); Berkenkamp et al. (2019); Ziomek et al. (2024) provided algorithms with theoretical guarantees when the kernel lengthscale is unknown. More recently, Ziomek et al. (2025) introduced an elimination-based algorithm with theoretical guarantees for an arbitrary set of discrete priors. Their algorithm, Prior-Elimination GP-UCB (PE-GP-UCB), selects the arm and prior which provide the most optimistic upper confidence bound (UCB). If a prior generates too many incorrect predictions, then it may be eliminated. The previous work has focused on optimistic UCB methods which are known to over-explore.
@@
In this work, we investigate the use of Thompson sampling for solving GP-bandit problems with unknown priors and we propose two algorithms. The first algorithm, Prior-Elimination GP-TS (PE-GP-TS), is an extension of PE-GP-UCB that replaces the doubly optimistic selection rule with posterior sampling and one less layer of optimism. We analyze the regret of PE-GP-TS. For the terms we can bound, we obtain a regret bound for PE-GP-TS of order ${\mathcal{O}}(\sqrt{T\log T|P|\hat{\gamma}_{T}})$ where $T$ is the horizon, $|P|$ is the number of priors and $\hat{\gamma}_{T}$ is the worst-case maximum information gain, which matches that of PE-GP-UCB. The second algorithm, HyperPrior GP-TS (HP-GP-TS), uses bi-level posterior sampling to efficiently explore the priors and arms. An UCB-based analysis of HP-GP-TS yields a regret bound of order ${\mathcal{O}}(\sqrt{T\log T\bar{\gamma}_{T}})$ (where $\bar{\gamma}_{T}$ is the average maximum information gain) plus a term that corresponds to the cost of learning the optimal prior. An information-theoretic analysis yields a regret bound of order ${\mathcal{O}}(\sqrt{T|{\mathcal{X}}|\log|{\mathcal{X}}|})$ where $|{\mathcal{X}}|$ is the number of arms.
We evaluate our methods on three sets of synthetic experiments and three experiments with real-world data. Across the experiments, our Thompson sampling based methods outperform PE-GP-UCB. Additionally, we find that the regret of HP-GP-TS does not increase with $|P|$ in our experiments. Finally, we analyze the priors selected by the algorithms and observe that HP-GP-TS selects the correct prior more often than the other algorithms.
@@
  • We theoretically analyze the regret of HP-GP-TS using a UCB framework and an information-theoretic framework which provides a regret bound of order ${\mathcal{O}}(\sqrt{T|{\mathcal{X}}|\log|{\mathcal{X}}|})$. Additionally, we analyze the regret of PE-GP-TS.
  • We experimentally evaluate our algorithms on both synthetic and real-world data, demonstrating that they achieve superior performance and that the regret of HP-GP-TS does not increase with $|P|$.
@@ 2 Background and problem statement
Figure 1: Agent deciding between priors based on observed data. The unknown function $f$ is sampled from a periodic prior and the observed data is in the top left. The agent discards the priors on the right since they do not match the new observation (star).
@@
We consider a sequential decision making problem where an agent repeatedly selects among a set of arms and receives a random reward whose mean depends on the selected arm and is unknown to the agent. The goal of the agent is to maximize the cumulative sum of rewards over a finite time horizon. We assume that the distribution of the means, the prior, is sampled from a set of priors, the hyperprior. An effective agent must distinguish which prior the means are sampled from to ensure it explores efficiently. We illustrate this in Fig.˜1.
Now, let us formally state the problem. Let ${\mathcal{X}}\subseteq[0,r]^{d}\subset{\mathbb{R}}^{d}$ denote the finite set of arms and $P$ a finite set of priors with associated prior mean and kernel functions $\mu_{1,p}:{\mathcal{X}}\mapsto{\mathbb{R}}$ and $k_{1,p}:{\mathcal{X}}\times{\mathcal{X}}\mapsto{\mathbb{R}}$, $\forall p\in P$. Let $p^{*}\in P$ denote the true prior and assume the expected reward function $f:{\mathcal{X}}\mapsto{\mathbb{R}}\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ is a sample from a Gaussian process with prior $p^{*}$. Both the function $f$ and the true prior $p^{*}$ are considered unknown. We will consider two settings: In the frequentist selection setting, the prior $p^{*}\in P$ is picked arbitrarily. In the Bayesian selection setting, the prior is sampled from a hyperprior $p^{*}\sim\mathcal{P}(P)$. To simplify notation, let $P_{1}$ denote the hyperprior.
Let $T$ denote the horizon. For time step $t=1,2,\ldots,T$, the agent selects an arm $x_{t}\in{\mathcal{X}}$ and observes the reward $y_{t}=f(x_{t})+\epsilon_{t}$ where $\{\epsilon_{t}\}_{t=1}^{T}$ are i.i.d. zero-mean Gaussian noise with variance $\sigma^{2}$. The goal of the agent is to select a sequence of arms $\{x_{t}\}_{t=1}^{T}$ that minimizes the regret $R(T)=\sum_{t\in[T]}f(x^{*})-f(x_{t})$ where $T]=\{1,\ldots,T\}$ and $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}f(x)$. In the Bayesian selection setting, we evaluate the agent based on the Bayesian regret ${\text{BR}}(T)={\mathbb{E}}\left[R(T)\right]$ where the expectation is taken over the prior $p^{*}$, the expected reward function $f$, the noise $\{\epsilon_{t}\}_{t=1}^{T}$ and the (potentially) stochastic selection of arms.
@@
A Gaussian process $f(x)\sim{\mathcal{GP}}(\mu,k)$ is a collection of random variables such that for any subset $\{x_{1},\ldots,x_{n}\}\subset{\mathcal{X}}$, the vector $f(x_{1}),\ldots,f(x_{n})]\in{\mathbb{R}}^{n}$ has a multivariate Gaussian distribution. The probabilistic nature of GPs make them very useful for defining and solving bandit problems where the arms are correlated. Given the history $H_{t}=\{(x_{i},y_{i})\}_{i=1}^{t-1}$, the posterior mean and kernel functions of a Gaussian process ${\mathcal{GP}}(\mu,k)$ are given by
@@
[…]
@@
The maximal information gain (MIG) is a measure of reduction in uncertainty of $f$ after observing the most informative data points up to a specified size. The MIG commonly occurs in regret bounds for GP bandit algorithms (Srinivas et al., 2012; Vakili et al., 2021) and its growth rate is strongly determined by the prior kernel of the GP. Hence, we will define the MIG for any fixed GP prior $p\in P$. Let ${\mathbf{y}}_{A}$ denote noisy observations of $f$ at the locations $A\subset{\mathcal{X}}$. Then, the MIG given prior $p\in P$, $\gamma_{T,p}$, is defined as
@@
where $I_{p}({\mathbf{y}}_{A};f)=H({\mathbf{y}}_{A}|p)-H({\mathbf{y}}_{A}|f,p)$ is the mutual information between ${\mathbf{y}}_{A}$ and $f$ given $p$, and $H(\cdot)$ denotes the entropy. To aid our analysis later, we also define the worst-case MIG as $\hat{\gamma}_{T}:=\max_{p\in P}\gamma_{T,p}$ and the average MIG as $\bar{\gamma}_{T}:={\mathbb{E}}_{p\sim P_{1}}[\gamma_{T,p}]$. For the RBF and Matérn kernels, $\gamma_{T,p}={\mathcal{O}}(\log^{d+1}(T))$ and $\gamma_{T,p}={\mathcal{O}}(T^{\frac{d}{2\nu+d}}\log^{\frac{2\nu}{2\nu+d}}(T))$ (Srinivas et al., 2012; Vakili et al., 2021).
@@
Plenty of previous work have proposed fully Bayesian approaches that integrate the acquisition function over the hyperposterior (Osborne et al., 2009; Benassi et al., 2011; Snoek et al., 2012; Hernández-Lobato et al., 2014; Wang & Jegelka, 2017; De Ath et al., 2021; Hvarfner et al., 2023). In contrast, HP-GP-TS optimizes a single hyperposterior sample instead of expected values over the hyperposterior.
Wang & de Freitas (2014) first derived regret bounds for GP bandits with unknown lengthscale for the Expected Improvement algorithm (Močkus, 1975). However, the proposed algorithm requires a lower bound on the lengthscale and the regret bound depends on the worst-case MIG. Later work by Berkenkamp et al. (2019) introduced Adaptive GP-UCB (A-GP-UCB) that continually lowers the lengthscale parameter. Given a sufficiently small lengthscale, the function $f$ lies within the reproducing kernel Hilbert space (RKHS) and the regular GP-UCB theory can be applied. However, A-GP-UCB lacks a stopping mechanism and will overexplore as the lengthscale continues to shrink. Recent work by Ziomek et al. (2025) introduced Prior-Elimination GP-UCB (PE-GP-UCB) for time-varying GP-bandits with unknown prior. Unlike the work before, the regret bound of PE-GP-UCB holds for arbitrary types of hyperparameters in the GP prior. PE-GP-UCB is doubly optimistic and selects the prior and arm with the highest upper confidence bound. PE-GP-UCB tracks the cumulative prediction error made by the selected priors and eliminates priors that exceed a threshold level.
Other works have introduced regret balancing algorithms that maintain a set of base learning algorithms and balance their selection frequency to achieve close to optimal regret (Abbasi-Yadkori et al., 2020; Pacchiano et al., 2020). Ziomek et al. (2024) built on this idea and introduced length-scale balancing GP-UCB which can adaptively explore smaller lengthscales but can return to longer ones, unlike A-GP-UCB.
The aforementioned works are based on UCB, EI, PI or regret balancing. However, another line of work has studied Thompson sampling in standard and linear bandits with unknown prior distribution (Kveton et al., 2021; Basu et al., 2021; Hong et al., 2022; Li et al., 2024). In their setting (meta or hierarchical bandits), the agent plays multiple bandit instances, either simultaneously or sequentially. The unknown means are sampled from the same (unknown) prior and by gathering knowledge across instances, the agent can solve later instances more efficiently once it has identified the prior. We emphasize that these methods have been studied only for standard stochastic and linear bandits, not for GP bandits.
Figure 2: a) Elimination procedure of PE-GP-TS. The solid lines correspond to posterior means and the shaded regions are confidence intervals. The figure has been adapted from Ziomek et al. (2025). The dashed lines are samples from the posteriors. b) Overview of HP-GP-TS. The orange star corresponds to $y_{t}$.